# Quantum chaos and the arrow of time

### v10

@@ 1 Introduction
All physical systems around us become more and more disordered as time progresses unless an outside agent spends energy and record information for the upkeep of the system. This phenomenon has a beautiful statistical explanation. There simply are too many possible disordered states compared to the number of possible ordered states. So, the system ends up most likely in a disordered state [1]. Entropy is the quantity which measures the extent of disorder. In terms of entropy, this general observation gives rise to the second law of thermodynamics. It states that the entropy of an isolated system does not decrease. This is the Planck statement of the second law of thermodynamics [2]. It gives rise to the thermodynamic arrow of time (called Time’s arrow in [3]). These results are familiar in the realm of classical physics. But our physical world is governed by quantum mechanics at microscopic level and so far a complete understanding of the quantum version of the second law is lacking. It is not clear if entropic consideration can fully explain second law. In quantum mechanics, wavefunction plays a central role. So, denser regions of the energy spectrum are not necessarily readily available. Transition amplitudes to the denser region of the energy spectrum also play a critical role. The central goal of this work is to study these transition amplitudes in quantum chaotic many-body systems.
@@
The second law of thermodynamics has been stated as different statements all of which are considered to be equivalent. We will work with the Kelvin statement of the second law. It states that no process is possible whose sole result is the complete conversion of heat into work [4, 5]. In this paper, we will be working with the microscopic theory so there will be no notion of heat. We will keep track of the energy of the system. So, the second law in the Kelvin form implies that it is impossible to convert energy from a single source into work without any other effect. In other word, if we perturb an otherwise closed system, the system cannot lose energy. If the system loses energy then the energy lost would have been completely converted into work. Examples of perturbing a system would be a momentary application of magnetic field in a spin system or shooting a pulse of laser at an electronic system. [^1] We are not turning on the magnetic field indefinitely in which case the Hamiltonian of the system would be changed forever. We are also not considering cases in which the system is coupled to another system. An example of coupling two systems would be connecting the piston of an expanding gas to some other mechanical system like a dynamo through a crankshaft to extract electrical energy. If we let the gas expand without connecting the piston to the crankshaft, the total energy of the gas system would remain constant. We have been considering positive temperature in our discussion. At negative temperature, the Kelvin statement implies that the system cannot gain energy when perturbed. At infinite temperature, there is no change in energy. In this paper, we will call the second law of thermodynamics in the form of the Kelvin statement as the arrow of time.
@@
In this work, we show that the arrow of time arises naturally in a isolated quantum chaotic system in the thermodynamic limit. We show that this result follows from an extension of Berry’s conjecture [6, 7] and equivalence of ensembles. Berry’s conjecture states that expectation values of observables in a single energy eigenstate of a quantum chaotic system are same as their microcanonical averages. Quantum chaos in many-body systems and the related question of thermalization of these systems have been areas of intense research in the last few decades [7]. Traditionally, quantum chaos is identified from the study of energy level statistics [8] and of delocalization measures for the complexity of eigenstates [9]. Berry’s conjecture deals with the diagonal elements of observables in the energy eigenstates. The off-diagonal matrix elements were first studied in detail in [10]. A formal ansatz for the matrix elements is now well known as Eigenstate thermalization hypothesis (ETH) [11, 12]. It is a comprehensive statement which explains how observables thermalize in a quantum system [13, 14, 7, 15]. So, it is the route to statistical mechanics for quantum systems. ETH has been shown to hold true for physically sensible observables of quantum chaotic systems while it does not hold true for quantum integrable systems. So for the rest of the paper, we will assume that quantum chaos and ETH imply each other. Berry’s conjecture is the diagonal part of ETH if we ignore the fluctuations. But ETH and thermalization do not immediately imply the second law of thermodynamics. We will discuss this matter in detail in section 2. So, the arrow of time gives rise to new constraints on the off-diagonal elements of ETH. The off-diagonal elements are the transition amplitudes between different energy eigenstates. The new constraint effectively says that, starting from an energy eigenstate, the bare transition amplitudes without the entropic factor to nearby energy eigenstates on the denser side of the spectrum should be equal to or greater than the bare transition amplitudes to the thinner side of the spectrum. The equality holds true in the thermodynamic limit when there is equivalence of ensembles. So, entropic consideration is not enough to give rise to the arrow of time. We will call these new constraints collectively as ETH-monotonicity. [^2] We thank the referee 3 for suggesting this terminology.
@@
It has previously been shown that the change in the diagonal entropy is non-negative under doubly stochastic evolution starting from a passive density matrix [16, 7]. This follows from the convex nature of the function $x\log x$. This result is true even for integrable systems. Thermal density matrix is an example of passive density matrix. But integrable systems do not thermalize after a finite perturbation. In the present work, we will be working with pure states which is the most appropriate approach for an isolated system. In the spirit of Berry’s conjecture and ETH, we will be working with individual energy eigenstates.
Just as we cannot keep track of the positions and momenta of all gas molecules in the classical theory of gas, we cannot keep track of the complex coefficients of the exponentially large number of energy eigenstates of a quantum system. So in quantum mechanics, we keep track of only certain observables which can be easily measured or calculated. These can be local operators or global operators but they are few-body operators. If we observe that these observables thermalize then we say that the system thermalizes. ETH is a criterion for thermalization of such observables so that their long time expectation values can be described by quantum statistical mechanics. Inevitably, the quantum operators corresponding to these observables are hermitian and non-fermionic in nature.
@@
$\langle m|\mathcal{O}|n\rangle\equiv\mathcal{O}_{mn}=\mathcal{O}(\bar{E})\delta_{mn}+e^{-S(\bar{E})/2}f(\bar{E},\omega)R_{mn}$ (2)
where $|m\rangle$ and $|n\rangle$ are energy eigenstates with energies $E_{m}$ and $E_{n}$ respectively and $\bar{E}=(E_{m}+E_{n})/2$, $\omega=E_{m}-E_{n}$. $S(\bar{E})$ is the entropy at energy $\bar{E}$. $\mathcal{O}(\bar{E})$ is equal to the expectation value of $\mathcal{O}$ in the microcanonical ensemble at energy $\bar{E}$ or other ensembles by the equivalence of ensembles. $\mathcal{O}(\bar{E})$ and $f(\bar{E},\omega)$ are smooth functions of their arguments. $f(\bar{E},\omega)$ is taken to be a real and positive function. $R_{mn}$ are pseudo-random variables with zero mean and unit variance [17]. $\mathcal{O}$ is hermitian and hence
$\mathcal{O}_{nm}=\mathcal{O}^{*}_{mn},\;R_{nm}=R^{*}_{mn},\;f(\bar{E},-\omega)=f(\bar{E},\omega)$ (3)
The $\omega$-dependence of the $f$-function has been a subject of study in many works [18, 19, 7, 20, 21, 22]. At large $|\omega|$, the $f$-function falls exponentially faster than $e^{-|\beta\omega|/4}$ where $\beta=1/T$ is the inverse effective temperature [19, 7, 20, 21]. Moreover, for small $\omega\sim 1/L^{2}$, where $L$ is the system size, $f$-function can be assumed to be a constant function as in case of random matrix theory (RMT) [18, 7, 20, 22].
@@
$f(S)=\check{f}(S(\bar{E}))=f(S^{-1}(S(\bar{E}))).$ (4)
We found that
@@
$\frac{f^{\prime}(\bar{E})}{f(\bar{E})}\;\xrightarrow{L\to\infty}\;0,\qquad L\;\frac{f^{\prime}(\bar{E})}{f(\bar{E})}\;=\;\text{a constant.}$. (5)
@@
R1 and R2 are analytic results. R3 and R4 are numerical results. We also provide some analytic motivations for R3. ETH-monotonicity comprises the results R2, R3 and R4. $\bar{E}$ is an extensive quantity but we will see that the flattening of $f(\bar{E})$ is not merely stretching of this function as the system size increases. Temperature $T$ is an intensive quantity so the result R3 will hold true even in the thermodynamic limit. Our numerical results are performed for finite systems with finite number of degrees of freedom. But we expect that our results will hold true even for continuum systems like quantum field theories.
We will be mostly working with pure states. The temperature $T$ of a pure state is the effective temperature calculated using equivalence of ensembles. Basically we calculate the temperature of the canonical ensemble having the same energy as the pure state. Correspondingly, the entropy (of even a single eigenstate) is the microcanonical entropy. For a clear perspective, the difference between canonical and microcanonical expectation values can be seen in [13]. In Appendix C, we will also find that the canonical entropy is always greater than the microcanonical entropy. The energy window of microcanonical ensemble $dE$ is usually considered to be an arbitrary constant as long as it does not grow extensively with system size. For this work, we need a precise definition of the energy window. We found a natural definition of the energy window by a comparative study of the microcanonical entropy and canonical entropy. Our choice of the energy window is such that the ratio of the canonical entropy and microcanonical entropy remains constant throughout the energy spectrum. The details can be found in Appendix C. We also found that this optimum energy window does not grow extensively with the system size. In fact, for the three different system sizes under consideration we found that the optimum energy window to be a constant and equal to $dE=0.8$. In Appendix C, we also compare our choice of the energy window with the energy window used in [13]. We believe that our new definition of the energy window is more natural since it arises from the comparison of entropies.
@@ 2 Thermalization does not imply the arrow of time
The most important consequence of ETH is thermalization of non-equilibrium excited states which are typical [19]. Typical states are states with sub-extensive energy variance. [^3] Note that this is the definition of a typical state in ETH works and it has no relation with other forms of quantum typicality studied in quantum statistical mechanics and quantum thermodynamics. If an observable satisfy the ETH statement then its expectation value in a typical state relaxes to the microcanonical average specified by the fixed conserved charges in the long time limit. Fluctuations away from the microcanonical average can be interpreted as thermal fluctuations [23], as long as we do not wait long enough to observe Poincare recurrence. This scenario is most popularly studied as quantum quenches where the non-equilibrium state is prepared by changing the Hamiltonian of the system [14, 15]. The initial Hamiltonian and the final Hamiltonian are different.
@@
One subtlety when preparing states using quantum quenches is that even integrable systems equilibrate to generalized Gibbs ensembles (GGEs) in the large system size limit, if the initial state has no long range correlation. Equilibration of quantum integrable systems after quantum quenches has been a topic of great interest in the last two decades [24, 25, 26, 14, 15, 27]. The non-equilibrium states are generalized typical states taking into account all the other conserved charges of the integrable system. This result follows from central limit theorem and the local nature of the Hamiltonians [13, 7]. The difference between chaotic systems and integrable systems arises when we consider finite perturbations for a finite duration of time (also called bump quenches in [28] and critical to critical quench in [26]). The initial Hamiltonian and the final Hamiltonian are same. A chaotic system still thermalizes after such finite perturbations [28] while an integrable system cannot be even described by a GGE in long time limit [26]. The temperature and other chemical potentials become imaginary. So in this manner, finite perturbations or bump quenches have a potential to bring out the stark contrast between non-equilibrium dynamics of quantum chaotic systems and quantum integrable systems.
@@
The apparent one-way evolution of a thermalization process, even when the microscopic dynamics is time-reversal invariant, has been considered to be an “arrow of time" [19]. But we note that thermalization is different from our definition of the thermodynamic arrow of time. First of all, just because a system thermalization does not imply that the change in energy $\Delta E$ has the same sign as the temperature $T$ which is what the Kelvin statement of the second law of thermodynamics implies. Indeed we expect that the system would thermalize back to a new thermal equilibrium after the perturbation has stopped, but this thermalization does not imply that $\Delta E$ will strictly be either positive or negative. In other word, although thermalization and the arrow of time under consideration are physical phenomena observed in a chaotic system, they do not imply each other.
@@ 3 Set-up and results
Consider a system with the time-independent Hamiltonian $H_{0}$ and initially in a state $|\psi\rangle$. The system is perturbed by a time-dependent term $\lambda(t)\mathcal{O}$ where $\mathcal{O}$ is a non-fermionic, hermitian operator. The total Hamiltonian is
$H(t)=H_{0}+\lambda(t)\mathcal{O}$ (6)
The perturbation is for a finite duration of time, say, from time $-d$ to $d$. The source $\lambda(t)$ is real and can be non-zero only for $-d<t<d$. This perturbing term is not necessarily small but it also cannot be arbitrarily big. The change in energy of the system is
$\Delta E=\langle\psi(t_{f})|H_{0}|\psi(t_{f})\rangle-\langle\psi(t_{i})|H_{0}|\psi(t_{i})\rangle$ (7)
@@
$\Delta E_{n}=\sum_{m}\Delta E_{nm}=\sum_{m}(E_{m}-E_{n})|\tilde{\lambda}(E_{m}-E_{n})|^{2}|\mathcal{O}_{mn}|^{2}+O(\lambda^{3})$ (8)
@@
For numerical calculations, we consider the XXZ spin chains with open boundary condition. With a large next-to-nearest neighbour interaction, the system is chaotic [29]. We consider a system of size $L$ with the number of up-spins $N$. So, total magnetization is $2N-L$. The numerical methods involved are described in Appendix E. The system is perturbed using the global spin-current operator or the kinetic energy operator (the hopping term). It has been shown that these operators satisfy the ETH statement in the chaotic XXZ spin chain [30]. Both the operators preserve the total magnetization.
$H_{0}=\sum_{i=1}^{L-1}\left[J_{xy}\left(S_{i}^{x}S_{i+1}^{x}+S_{i}^{y}S_{i+1}^{y}\right)+J_{z}S_{i}^{z}S_{i+1}^{z}\right]+\sum_{i=1}^{L-2}J^{\prime}_{z}S_{i}^{z}S_{i+2}^{z}$ (9)
$\mathcal{O}_{SC}=\sum_{i=1}^{L-1}J_{xy}\left(S_{i}^{x}S_{i+1}^{y}-S_{i}^{y}S_{i+1}^{x}\right),\qquad\mathcal{O}_{KE}=\sum_{i=1}^{L-1}J_{xy}\left(S_{i}^{x}S_{i+1}^{x}+S_{i}^{y}S_{i+1}^{y}\right).$ (10)

### v12

@@ 1 Introduction
All physical systems around us become more and more disordered as time progresses unless an outside agent spends energy and record information for the upkeep of the system. This phenomenon has a beautiful statistical explanation. There simply are too many possible disordered states compared to the number of possible ordered states. So, the system ends up most likely in a disordered state Greiner1995 . Entropy is the quantity which measures the extent of disorder. In terms of entropy, this general observation gives rise to the second law of thermodynamics. It states that the entropy of an isolated system does not decrease. This is the Planck statement of the second law of thermodynamics planck1903treatise . It gives rise to the thermodynamic arrow of time (called Time’s arrow in Eddington1927 ). These results are familiar in the realm of classical physics. But our physical world is governed by quantum mechanics at microscopic level and so far a complete understanding of the quantum version of the second law is lacking. It is not clear if entropic consideration can fully explain the second law. In quantum mechanics, wavefunction plays a central role. So, denser regions of the energy spectrum are not necessarily readily available. Transition amplitudes to the denser region of the energy spectrum also play a critical role. For example, if the transition amplitudes to the energy eigenstates in the denser region is infinitesimally small then the wavefunction of the system would not transit towards this denser region of the energy spectrum. The central goal of this work is to study these transition amplitudes in quantum chaotic many-body systems.
@@
The second law of thermodynamics has been stated as different statements all of which are considered to be equivalent. We will work with the Kelvin statement of the second law. It states that no process is possible whose sole result is the complete conversion of heat into work Kardarbook2007 ; huang2008statistical . In this paper, we will be working with the microscopic theory so there will be no notion of heat. We will keep track of the energy of the system. So, the second law in the Kelvin form implies that it is impossible to convert energy from a single source into work without any other effect. In other word, if we perturb an otherwise closed system, the system cannot lose energy. If the system loses energy then the energy lost would have been completely converted into work. Examples of perturbing a system would be a momentary application of magnetic field in a spin system or shooting a pulse of laser at an electronic system. [^1] We are not turning on the magnetic field indefinitely in which case the Hamiltonian of the system would be changed forever. We are also not considering cases in which the system is coupled to another system. An example of coupling two systems would be connecting the piston of an expanding gas to some other mechanical system like a dynamo through a crankshaft to extract electrical energy. If we let the gas expand without connecting the piston to the crankshaft, the total energy of the gas system would remain constant. We have been considering positive temperature in our discussion. At negative temperature, the Kelvin statement implies that the system cannot gain energy when perturbed. At infinite temperature, there is no change in energy. In this paper, we will call the second law of thermodynamics in the form of the Kelvin statement as the arrow of time.
@@
In this work, we show that the arrow of time arises naturally in a isolated quantum chaotic system in the thermodynamic limit. We show that this result follows from an extension of Berry’s conjecture Berry_1977 ; DAlessio:2015qtq and equivalence of ensembles. Berry’s conjecture states that expectation values of observables in a single energy eigenstate of a quantum chaotic system are same as their microcanonical averages. Quantum chaos in many-body systems and the related question of thermalization of these systems have been areas of intense research in the last few decades DAlessio:2015qtq . Traditionally, quantum chaos is identified from the study of energy level statistics Bohigas:1983er and of delocalization measures for the complexity of eigenstates Santos_PRE_2010 . Berry’s conjecture deals with the diagonal elements of observables in the energy eigenstates. The off-diagonal matrix elements were first studied in detail in PhysRevA.34.591 . For many-body systems, a formal ansatz for the matrix elements is now well known as eigenstate thermalization hypothesis (ETH) PhysRevA.43.2046 ; Srednicki:1994mfb . It is a comprehensive statement which explains how observables thermalize in a quantum system Rigol_2008nature ; Polkovnikov:2010yn ; DAlessio:2015qtq ; Gogolin:2016hwy . So, it is the route to statistical mechanics for quantum systems. ETH has been shown to hold true for physically sensible observables of quantum chaotic systems while it does not hold true for quantum integrable systems. So for the rest of the paper, we will assume that quantum chaos and ETH imply each other. Berry’s conjecture is the diagonal part of ETH if we ignore the fluctuations. But ETH and thermalization do not immediately imply the second law of thermodynamics. We will discuss this matter in detail in section 2. So, the arrow of time gives rise to new constraints on the off-diagonal elements of ETH. The off-diagonal elements are the transition amplitudes between different energy eigenstates. The new constraint effectively says that, starting from an energy eigenstate, the bare transition amplitudes without the entropic factor to nearby energy eigenstates on the denser side of the spectrum should be equal to or greater than the bare transition amplitudes to the thinner side of the spectrum. The equality holds true in the thermodynamic limit when there is equivalence of ensembles. So, entropic consideration is not enough to give rise to the arrow of time. We will call these new constraints collectively as ETH-monotonicity.
@@
It has previously been shown that the change in the diagonal entropy is non-negative under doubly stochastic evolution starting from a passive density matrix Polkovnikov_2011 ; DAlessio:2015qtq . This follows from the convex nature of the function $x\log x$. This result is true even for integrable systems. Thermal density matrix is an example of passive density matrix. But integrable systems do not thermalize after a finite perturbation. In the present work, we will be working with pure states which is the most appropriate approach for an isolated system. In the spirit of Berry’s conjecture and ETH, we will be working with individual energy eigenstates.
Just as we cannot keep track of the positions and momenta of all gas molecules in the classical theory of gas, we cannot keep track of the complex coefficients of the exponentially large number of energy eigenstates of a quantum system. So in quantum mechanics, we keep track of only certain observables which can be easily measured or calculated. These can be intensive (local) operators or extensive (global) operators but they are few-body operators. If we observe that these observables thermalize then we say that the system thermalizes. ETH is a criterion for thermalization of such observables so that their long time expectation values can be described by quantum statistical mechanics. Inevitably, the quantum operators corresponding to these observables are hermitian and non-fermionic in nature.
@@
$\langle m|\mathcal{O}|n\rangle\equiv\mathcal{O}_{mn}=\mathcal{O}(\bar{E})\delta_{mn}+e^{-S(\bar{E})/2}f(\bar{E},\omega)R_{mn}\$(1.2)
where $|m\rangle$ and $|n\rangle$ are energy eigenstates with energies $E_{m}$ and $E_{n}$ respectively and $\bar{E}=(E_{m}+E_{n})/2$, $\omega=E_{m}-E_{n}$. $S(\bar{E})$ is the entropy at energy $\bar{E}$. $\mathcal{O}(\bar{E})$ is equal to the expectation value of $\mathcal{O}$ in the microcanonical ensemble at energy $\bar{E}$ or other ensembles by the equivalence of ensembles. $\mathcal{O}(\bar{E})$ and $f(\bar{E},\omega)$ are smooth functions of their arguments. $f(\bar{E},\omega)$ is taken to be a real and positive function. $R_{mn}$ are pseudo-random variables with zero mean and unit variance Foini:2018sdb . $\mathcal{O}$ is hermitian and hence
$\mathcal{O}_{nm}=\mathcal{O}^{*}_{mn},\;R_{nm}=R^{*}_{mn},\;f(\bar{E},-\omega)=f(\bar{E},\omega)$ (1.3)
The $\omega$-dependence of the $f$-function has been a subject of study in many works Khatami_2013 ; Srednicki_1999 ; DAlessio:2015qtq ; Dymarsky:2018ccu ; Murthy:2019fgs ; Wang:2021mtp . At large $|\omega|$, the $f$-function falls exponentially faster than $e^{-|\beta\omega|/4}$ where $\beta=1/T$ is the inverse effective temperature Srednicki_1999 ; DAlessio:2015qtq ; Dymarsky:2018ccu ; Murthy:2019fgs . Moreover, for small $\omega\sim 1/L^{2}$, where $L$ is the system size, $f$-function can be assumed to be a constant function as in case of random matrix theory (RMT) Khatami_2013 ; DAlessio:2015qtq ; Dymarsky:2018ccu ; Wang:2021mtp .
@@
$f(S)=\check{f}(S(\bar{E}))=f(S^{-1}(S(\bar{E}))).\$(1.4)
$S^{-1}$ function can be multivalued. We have to choose the appropriate value by crosschecking with values of $T$ or $\beta$. We found that
@@
$\frac{f^{\prime}(\bar{E})}{f(\bar{E})}\;\xrightarrow{L\to\infty}\;0,\qquad L\;\frac{f^{\prime}(\bar{E})}{f(\bar{E})}\;=\;\text{a constant.}$ (1.5)
@@
R1 and R2 are analytic results. R3 and R4 are numerical results. We also provide some analytic motivations for R3. ETH-monotonicity comprises the results R2, R3 and R4. $\bar{E}$ is an extensive quantity but we will see that the flattening of $f(\bar{E})$ is not merely stretching of this function as the system size increases. Temperature $T$ is an intensive quantity so the monotonically increasing property of $f(T)$ will hold true even in the thermodynamic limit. Our numerical results are performed for finite systems with finite number of degrees of freedom. But we expect that our results will hold true even for continuum systems like quantum field theories.
We will be mostly working with pure states. The temperature $T$ of a pure state is the effective temperature calculated using equivalence of ensembles. Basically we calculate the temperature of the canonical ensemble having the same energy as the pure state. Correspondingly, the entropy (of even a single eigenstate) is the microcanonical entropy. For a clear perspective, the difference between canonical and microcanonical expectation values can be seen in Rigol_2008nature . In Appendix C, we will also find that the canonical entropy is always greater than the microcanonical entropy. The energy window of microcanonical ensemble $dE$ is usually considered to be an arbitrary constant as long as it does not grow extensively with system size. For this work, we need a precise definition of the energy window. We found a natural definition of the energy window by a comparative study of the microcanonical entropy and canonical entropy. Our choice of the energy window is such that the ratio of the canonical entropy and microcanonical entropy remains constant throughout the energy spectrum. The details can be found in Appendix C. We also found that this optimum energy window does not grow extensively with the system size. In fact, for the three different system sizes under consideration we found that the optimum energy window to be a constant and equal to $dE=0.8$. In Appendix C, we also compare our choice of the energy window with the energy window used in Rigol_2008nature . We believe that our new definition of the energy window is more natural since it arises from the comparison of entropies.
@@ 2 Thermalization does not imply the arrow of time
The most important consequence of ETH is thermalization of non-equilibrium excited states which are typical Srednicki_1999 . Typical states are states with sub-extensive energy variance. [^2] Note that this is the definition of a typical state in ETH works and it has no relation with other forms of quantum typicality studied in quantum statistical mechanics and quantum thermodynamics. If an observable satisfy the ETH statement then its expectation value in a typical state relaxes to the microcanonical average specified by the fixed conserved charges in the long time limit. Fluctuations away from the microcanonical average can be interpreted as thermal fluctuations Srednicki:1995pt , as long as we do not wait long enough to observe Poincare recurrence. This scenario is most popularly studied as quantum quenches where the non-equilibrium state is prepared by changing the Hamiltonian of the system Polkovnikov:2010yn ; Gogolin:2016hwy . The initial Hamiltonian and the final Hamiltonian are different.
@@
One subtlety when preparing states using quantum quenches is that even integrable systems equilibrate to generalized Gibbs ensembles (GGEs) in the large system size limit, if the initial state has no long range correlation. Equilibration of quantum integrable systems after quantum quenches has been a topic of great interest in the last two decades Rigol_2007 ; Mandal:2015jla ; Mandal:2015kxi ; Polkovnikov:2010yn ; Gogolin:2016hwy ; Banerjee:2019ilw . The non-equilibrium states are generalized typical states taking into account all the other conserved charges of the integrable system. This result follows from central limit theorem and the local nature of the Hamiltonians Rigol_2008nature ; DAlessio:2015qtq . The difference between chaotic systems and integrable systems arises when we consider finite perturbations for a finite duration of time (also called critical to critical quench in Mandal:2015kxi and bump quenches in Bhattacharya:2018fkq ). The initial Hamiltonian and the final Hamiltonian are same. A chaotic system still thermalizes after such finite perturbations Bhattacharya:2018fkq while an integrable system cannot be even described by a GGE in long time limit Mandal:2015kxi . The temperature and other chemical potentials become imaginary. So in this manner, finite perturbations or bump quenches have a potential to bring out the stark contrast between non-equilibrium dynamics of quantum chaotic systems and quantum integrable systems.
@@
The apparent one-way evolution of a thermalization process, even when the microscopic dynamics is time-reversal invariant, has been considered to be an “arrow of time” Srednicki_1999 . But we note that thermalization is different from our definition of the thermodynamic arrow of time. First of all, just because a system thermalization does not imply that the change in energy $\Delta E$ has the same sign as the temperature $T$ which is what the Kelvin statement of the second law of thermodynamics implies. Indeed we expect that the system would thermalize back to a new thermal equilibrium after the perturbation has stopped, but this thermalization does not imply that $\Delta E$ will strictly be either positive or negative. In other word, although thermalization and the arrow of time under consideration are physical phenomena observed in a chaotic system, they do not imply each other.
@@ 3 Set-up and results
Consider a system with the time-independent Hamiltonian $H_{0}$ and initially in a state $|\psi\rangle$. The system is perturbed by a time-dependent term $\lambda(t)\mathcal{O}$ where $\mathcal{O}$ is a non-fermionic, hermitian operator. $\mathcal{O}$ could be either an intensive operator or an extensive operator. The total Hamiltonian is
$H(t)=H_{0}+\lambda(t)\mathcal{O}$ (3.1)
The perturbation is for a finite duration of time, say, from time d-d to $d$. The source $\lambda(t)$ is real and can be non-zero only for d<t<d-d<t<d. This perturbing term is not necessarily small but it also cannot be arbitrarily big. The change in energy of the system is
$\Delta E=\langle\psi(t_{f})|H_{0}|\psi(t_{f})\rangle-\langle\psi(t_{i})|H_{0}|\psi(t_{i})\rangle\$(3.2)
@@
$\Delta E_{n}=\sum_{m}\Delta E_{nm}=\sum_{m}(E_{m}-E_{n})|\tilde{\lambda}(E_{m}-E_{n})|^{2}|\mathcal{O}_{mn}|^{2}+O(\lambda^{3})\$(3.3)
@@
For numerical calculations, we consider the XXZ spin chains with open boundary condition. With a large next-to-nearest neighbour interaction, the system is chaotic Gubin_2012 . We consider a system of size $L$ with the number of up-spins $N$. So, total magnetization is 2NL2N-L. The numerical methods involved are described in Appendix E. The system is perturbed using the global spin-current operator or the kinetic energy operator (the hopping term). It has been shown that these operators satisfy the ETH statement in the chaotic XXZ spin chain Steinigeweg_2013 . Both the operators preserve the total magnetization.
$H_{0}=\sum_{i=1}^{L-1}\left[J_{xy}\left(S_{i}^{x}S_{i+1}^{x}+S_{i}^{y}S_{i+1}^{y}\right)+J_{z}S_{i}^{z}S_{i+1}^{z}\right]+\sum_{i=1}^{L-2}J^{\prime}_{z}S_{i}^{z}S_{i+2}^{z}$ (3.4)
$\mathcal{O}_{SC}=\sum_{i=1}^{L-1}J_{xy}\left(S_{i}^{x}S_{i+1}^{y}-S_{i}^{y}S_{i+1}^{x}\right),\qquad\mathcal{O}_{KE}=\sum_{i=1}^{L-1}J_{xy}\left(S_{i}^{x}S_{i+1}^{x}+S_{i}^{y}S_{i+1}^{y}\right).\$(3.5)